# Model Editing at Scale leads to Gradual and Catastrophic Forgetting

### 15 Jan 2024 v1

@@ 3.3 Downstream Evaluation of Edited Models
One implicit feature expected out of all model editing methods is that as a fact is edited or inserted into the model memory, it should not effect the model’s ability to perform its regular functions. This means that knowledge editing should not effect the model’s ability to perform common NLP tasks which the model is used for. We call this an implicit assumption because to the best of our knowledge, none of the previous works try to directly measure the effect of model editing on downstream tasks. Meng et al. (2022a, b) try to capture this effect in some form by calculating the generation entropy, which measures weighted average of the bi-gram and tri-gram entropy of text generated by the post-edit model. The idea is to measure model degradation when generating text using a post-edit model. Ideally, we want the generation entropy of cases with lots of repetitions (a common failure case of model editing (Meng et al., 2022a, b)) to be lower than coherently generated text. We find that this metric is unable to directly quantify the degradation of model ability as multiple edits are made to the model. We explain this in more detail and give examples of failures of generation fluency metric in appendix A.1.
@@
We take a more direct approach to quantify model degradation by measuring the performance of the post-edit model on common downstream NLP tasks. We chose two tasks from the popular GLUE benchmark Wang et al. (2018) - sentiment analysis (SST2) Socher et al. (2013) and paraphrase detection (MRPC) Dolan and Brockett (2005) for this. Both are binary classification tasks and we use a balanced subset of 200 test examples and evaluate the models using the accuracy metric. We evaluate the downstream model performance on these tasks as a function of number of edits made to the model in a few-shot setting. The in-context examples and prompt template used for evaluation are presented in Tables 2 and 3.
@@
Figure 3 depicts the effect of model editing on downstream tasks as a function of the number of edits made to the model. We again see two regions of model degradation. Initially, there is a gradual decline in downstream performance of the model for both SST-2 and MRPC tasks with an increasing number of edits made to the model. We then see an inflection point with an abrupt loss of ability of the model to perform either tasks, which coincides with the point of sudden decrease in editability of the model (section 3.1) and a catastrophic increase in forgetting previously edited facts (section 3.2). Additionally, we want to point the user to Sample 1 in Figure 3 which contains the maximum number of edits made to the model before the inflection point. We see that the model is no longer able to perform paraphrase detection long before the inflection point, showing that as the model is edited multiple times, it continuously loses its ability to perform downstream functions and may be unusable for certain downstream tasks even before the inflection point.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 4: This figure shows the normalized L2 distance between the edited layer weights and its original weights as a function of the number of edits made to the model. The editing algorithm is ROME and the model being edited is GPT2-XL.
@@
While the inflection point is a big concern for model editing methods, there is also a gradual decrease of general ability of the model even prior to that point, which is only visible if the model is evaluated on a downstream task. This shows the usefulness of evaluation of model editing methods on downstream tasks, which we urge the research community to adopt along with other knowledge editing metrics. We define the first region where the model progressively loses its ability to do prior tasks (like recalling previously edited facts or performing downstream tasks) as gradual forgetting, juxtaposing it with catastrophic forgetting. Note that forgetting here does not just refer to forgetting previously edited facts but a general loss of ability to perform a certain function. We find that sequential editing of a model always leads to these two phases of forgetting - gradual forgetting and catastrophic forgetting.
@@ 3.4 The Source of Forgetting
So far, we’ve seen that sequentially editing of multiple facts in LLMs leads to the model gradually forgetting previously edited facts and losing the ability to be useful for downstream tasks. This is followed by an abrupt inflection point, which not only leads to a forgetting almost all previously edited facts, but also a complete loss of model ability to perform regular NLP tasks, thus rendering the model useless. This inflection point is a fundamental feature of ROME which can be seen across all samples of multiple sequential edits, and is a realization of extreme catastrophic forgetting. But is this point an outcome of continuous editing of the model or the result of a single edit to the model? What is the reason behind these two phases of forgetting? In this section, we answer these question in more detail.
Model editing methods are designed with the objective of editing or inserting specific facts stored inside the model without changing all the weights of the model. A precursor to be able to do this is localizing a fact down to specific neurons inside a model and then only changing the weights of the identified neurons. The ROME method is built on the assumption that a fact can be changed by changing the weights of any one out of a set of knowledge-storing layers of a model while keeping the rest of the model the same, which is showed to work empirically and backed by causal tracing experiments Meng et al. (2022a). Each time we make such edits, the edited layer becomes slightly different from its original version. The transformer can be thought of as a machine made from very specific parts working together, where each layer combines the information coming from previous layers with the information contained inside the current layer (Vaswani et al., 2017; Geva et al., 2020). To be able to do this, each layer must be able to understand the signal coming from the previous layers. In simpler words, there is a notion of compatibility between the different layers of the transformer when they are trained together. As we edit one specific layer of the model continuously while keeping the rest of the model constant, we are constantly changing one part of the model while keeping the remaining part the same. Such a procedure is bound to reach a point where the layer that is changed becomes so different from its original version that this compatibility is destroyed and the other parts of the transformer are unable to makes sense of the incoming signal from the layer being edited.
@@
Figure 5: This figure shows the editing proficiency of MEMIT on GPT-2XL for Sample 1 over 5000 sequential edits made to the model.
@@
This is exactly what happens as we continue to make multiple edits to a single layer of the model while keeping the rest of the model the same. This can be seen in Figure 4. Figure 4 shows the normalized [^5] We first take the L2 norm between original and post-edit weights of the layer being edited, and then normalize it by number of neurons in the layer. L2 distance between the weights of the edited layer and the original weights of the layer as a function the number of edits made to the model. We find that as more edits are made to the model, the distance between the original and edited layer continuously increases until it suddenly explodes. This is the point where the edited layer becomes incompatible with the rest of the model. At this point, the model breaks down and catastrophically forgets previously learnt facts, loses its ability to do downstream tasks and its ability to be correct by model editing methods. The gradual increase in distance between the original weight and new weights leads to the gradual region of forgetting, whereas the spike in the distance with a single updates leads to catastrophic forgetting.
We can clearly see that the spikes in Figure 4 exactly coincide with the spikes in Figures 2 and Figure 3. This also explains the multiple spikes in Figure 2. As the model reaches the inflection point, the model editability also goes down but ROME is still able to edit some facts with limited accuracy. At this point, the model is over-fitting on the edited facts as it is only able to retrieve the recently edited facts and has completely forgotten previously edited knowledge and the ability to do downstream tasks. But as the model is edited further, we again see a spike in distance from the weights of the model at the previous spike, and the model again goes through catastrophic forgetting of the recently learnt facts.
@@
Finally, we take a deeper look at the specific edits that cause the inflection point. We call these edits disabling edits, as they disable the model and make it unusable for downstream tasks. Are these disabling edits a result of continuous sequential editing that accumulates over time or of one specific fact that is especially hard for a model to learn? We find that when we edit the facts corresponding to the disabling edits as the first edit made to the model, the model is still left completely disabled, that is, it is still unable to perform any downstream tasks, and the normalized distance of the layer weights from the original weights is comparable to the distances seen around the spikes. This shows that the disabling edits in ROME are not a result of continuous sequential editing of the model, but are a fundamental limitation of ROME. We can describe disabling edits as facts that ROME is unable to successfully edit without crippling the model. This also disproves the fundamental hypothesis behind ROME that knowledge in LLMs can be edited by changing the weights of a single layer inside the model.
We also want to point out that while existence of such disabling edits are a fundamental limitation of ROME, it can also be used as a feature. If we can test the editability of a fact prior to editing by checking if it is a disabling fact, we can possibly prevent catastrophic forgetting when using ROME. Such explorations are beyond the scope of this paper and should be part of future research.
@@ 4 Scaling MEMIT
In this section, we evaluate the performance of the Mass-Editing Memory in a Transformer (MEMIT) method Meng et al. (2022b) when multiple sequential edits are made to the same model. We will follow the same procedure as followed in section 3, first evaluating the editing proficiency, gradual and catastrophic forgetting and loss of performance on downstream tasks. We perform sequential editing on GPT2-XL using a random subset of 5000 examples from the CounterFact dataset when using MEMIT. This subset is a continuation of sample 1 in section 3. The behavior of MEMIT when sequentially editing GPT2-XL is shown in Figure 5.
@@
Figure 4(a) shows the editing proficiency of MEMIT as a function of the number of edits made to the model. Note that here we edit one fact at a time for MEMIT. While MEMIT is able to make batched edits, we leave that analysis for future work. The dotted lines show a window size of 5 previous edits, whereas solid lines show a window size of 50 previous edits, same as in Figure 1. We see that the efficacy score for MEMIT is not as high as ROME. This means that knowledge edits made via MEMIT are not always as successful as in ROME. We also see a continuous decline of neighborhood score for MEMIT as seen for ROME, showing that editing facts possibly start affecting other facts stored in the model.
Figure 4(b) shows the percentage of successfully edited facts that get forgotten as new facts are edited using MEMIT. We again begin to see two phases of forgetting - gradual and catastrophic. Forgetting is evaluated in Figure 4(b) after every 500 edits beyond 2000 edits for computational reasons, which is why the plot looks spiky afterwards. For this reason, we want to clearly point out the region of gradual forgetting to the reader, which goes on until approximately 4000 edits, which is followed by the region of catastrophic forgetting after which the model forgets all previously learnt facts. When compared with the gradual forgetting region of Figure 2, we find that MEMIT forgets fewer previously edited facts when compared to ROME. This is seen across multiple samples of edits made using MEMIT, although we show only one due to space constraints. As a comparison, model edited with ROME forgets approximately 30% facts after 600 edits in sample 1 (Figure 1(a)), whereas when editing the same sample with examples in the exact same order, MEMIT forgets approximately 10% previously edited facts. Additionally, we don’t encounter an inflection point for a larger number of edits when compared to ROME. This is true for multiple samples edited using MEMIT. This shows that although MEMIT is prone to gradual forgetting, it has a longer gradual forgetting phase, thus making it more stable when making a larger number of sequential edits to the model.
Figure 6: This figure shows the normalized L2 distance between the edited layer weights and its original weights as a function of the number of edits made to the model. The editing algorithm is MEMIT and the model being edited is GPT2-XL.
@@
Finally, the effect of the number of edits on the downstream performance of the model can be seen in Figure 4(c). We see that the model loses its ability to do tasks gradually as more number of edits are made to the model. We again see two regions of forgetting - a gradual loss in ability and then a sudden drop in the model’s ability to do downstream tasks. The model loses the ability to do paraphrase detection long before the point of catastrophic forgetting, thus showcasing that the model can lose its ability of performing certain downstream tasks even before a disabling edit cripples the model.
@@
Figure 6 shows the distance of the edited layers from the respective original layers. We find that the distance from respective original layer increases gradually until 4000 edits, which is also the region of gradual forgetting of previously edited facts and downstream ability. After this, we find a sudden spike in the distance between the edited layers and the original layers, which coincides with the points of catastrophic forgetting as seen in previous plots. We also evaluate if disabling edits are fundamental to MEMIT, and find that if the fact that disables the model after a sequence of edit is edited first in the model, it does not lead to catastrophic forgetting. This means that in MEMIT, the effect of disabling edits is normalized. This supports the hypothesis behind MEMIT that facts are possibly stored across multiple layers and to be able to successfully edit facts, we need to edit the weights across each of these fact storing layers.
5 Limitations of ROME and MEMIT at Scale
We find that ROME and MEMIT have significant functional differences. ROME has almost 100% efficacy when editing facts although the model ends up forgetting a larger number of previously edited facts as multiple edits are made sequentially. MEMIT on the other hand is unable to successfully edit all facts, but fewer facts get forgotten as multiple edits are made to the model. Both ROME and MEMIT have low specificity and edited facts consistently bleed into other knowledge stored inside the model, especially with larger number of edits made to the model. MEMIT is also more robust to a single de-stabilizing edit when compared to ROME. We conjecture this is because a fact is stored within multiple layers of a model, and editing the weights of a single layer to edit facts can lead to instability in the model. Additionally, we see that both ROME and MEMIT are prone to forgetting in a machine learning sense, losing the ability to recall previously edited facts, to perform downstream function and the ability to be successfuly edited by these algorithms. This makes these algorithms hard to scale and not usable for practical model editing purposes.
@@ 6 Conclusion
In this paper we analyze two popular model editing techniques - ROME and MEMIT, at scale. We use these methods to make multiple sequential edits to the same model and find that they fail in multiple ways. We discover that the model being edited with these techniques undergoes forgetting of previous knowledge and skills in two phases. Initially, it gradually forgets previously edited facts and loses the ability to do downstream tasks, a phase which we call gradual forgetting. After that, the model abruptly loses all coherence and function including the ability to recall previously edited facts and perform downstream tasks, which is a realization of catastrophic forgetting. We also find that the source of these two phases of forgetting is that the layers being edited with these methods slowly drift away from their original weight values, thus becoming incompatible with the rest of the model.
Practical use of model editing requires us to be able to make multiple sequential edits to a model. The above findings indicate serious limitations of current state-of-the-art model editing methods at scale. We find that these model editing methods, like other fine tuning techniques, are prone to catastrophic forgetting. To be able to scale such methods, we not only need to have high efficacy, specificity and generalization, but we also need these methods to preserve the model’s existing abilities. Thus, it becomes essential to evaluate post-edit models on downstream tasks. Also, as multiple facts are edited inside a model, the model should remember previously edited facts. With this paper, we call for an improved evaluation of model editing techniques, including evaluating model performance on downstream tasks and ability to recall previously edited facts.
Finally, we want to stress upon the implications of the two phases of forgetting discovered in this paper. Gradual forgetting makes model editing techniques become increasingly less effective as we sequentially edit facts, and hence limits their usefulness at scale. While catastrophic forgetting, which renders the model practically useless, caps the extent to which we can scale these methods. Thus we need to create model editing techniques that can counteract both gradual forgetting and catastrophic forgetting when scaled.

### 20 Feb 2024 v2

@@ 3.3 Downstream Evaluation of Edited Models
One implicit feature expected out of all model editing methods is that as a fact is edited or inserted into model memory, it does not affect the model’s ability to perform its regular functions. This means that knowledge editing should not affect the model’s ability to perform common NLP tasks which the model is used for. We call this an implicit assumption because to the best of our knowledge, none of prior works try to directly measure the effect of model editing on downstream tasks [^4] A concurrent work also proposes evaluating model editing on downstream tasks Gu et al. (2024), which is completely coincidental. While their work solely focus on downstream evaluation, our work goes beyond that..
@@
We quantify model degradation by measuring the performance of the post-edit model on common downstream NLP tasks. We choose four tasks from the popular GLUE benchmark Wang et al. (2018) - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation. All tasks are binary classification tasks and we use a balanced subset of 200 test examples and evaluate the models using the F1-metric every few edits. While a more comprehensive selection of downstream tasks can be created, in this paper, our aim is to show the importance of such an evaluation at scale. We leave a more exhaustive analysis of model editing methods on downstream tasks for future work. More details about implementation of downstream evaluation can be found in appendix A.4.
@@
Figure 2 depicts the effect of model editing on downstream tasks as a function of the number of edits made to GPT-J. We see consistent model degradation as edits are made to the model across different methods of editing. There is a gradual but continuous degradation of model performance using FT-C, while the effect is sudden for MEND. For ROME, we again see two regions of model degradation. Initially, there is a gradual decline in downstream performance of the model with an increasing number of edits made to the model. We then see an inflection point with an abrupt loss of ability of the model to perform any downstream task, which coincides with the point of sudden decrease in editability of the model (Figure 0(c)) and a catastrophic increase in forgetting previously edited facts (Figure 0(d)).
@@
While the inflection point is a big concern for model editing methods, there is also a gradual decrease of general ability of the model even prior to that point, which is only visible if the model is evaluated on downstream tasks. This shows the usefulness of evaluation of model editing methods on downstream tasks, which we urge the research community to adopt along with other knowledge editing metrics. We define the first region where the model progressively loses its ability to do prior tasks (like recalling previously edited facts or performing downstream tasks) as gradual forgetting, juxtaposing it with catastrophic forgetting. Note that forgetting here does not just refer to forgetting previously edited facts but a general loss of ability to perform a certain function. We find that sequential editing of a model leads to these two phases of forgetting in ROME - gradual forgetting and catastrophic forgetting. We associate the region beyond the point of catastrophic forgetting with catastrophic forgetting as, after this point, model editing becomes ineffective and the model is almost unusable. For FT-C, we only observe a gradual forgetting, whereas For ROME, we see both gradual and catastrophic forgetting.
DISABLING EDITS NORMAL EDITS
DISTANCE
$3.339\times 10^{-4}$
$8.156\times 10^{-7}$
Table 2: Table showing average distance between edited layer weights from its original weights for disabling versus normal edits.
@@ 3.4 The Source of Forgetting
So far, we’ve seen that sequential editing of multiple facts in LLMs leads to the model gradually forgetting previously edited facts and losing the ability to be useful for downstream tasks. For ROME, this is followed by an abrupt inflection point, which not only leads to forgetting almost all previously edited facts, but also a complete loss of model ability to perform regular NLP tasks, thus rendering the model useless. Generation examples of model at this point can be seen in Table 8. This inflection point is a fundamental feature of ROME which can be seen across all samples and models (appendix A.4), and is a realization of extreme catastrophic forgetting. But is this point an outcome of continuous editing of the model or the result of a single edit to the model? What is the reason behind these two phases of forgetting? In this section, we answer these question in more detail.
Model editing methods are designed with the objective of editing or inserting specific facts stored inside the model without changing all the weights of the model (Dai et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Yao et al., 2023). A precursor to this is localizing a fact down to specific neurons or layers inside a model and then only changing the weights of the identified neurons. The ROME method is built on the assumption that a fact can be changed by changing the weights of any one out of a set of knowledge-storing layers of a model while keeping the rest of the model the same, which is showed to work empirically and backed by causal tracing experiments Meng et al. (2022a). Each time we make such edits, the edited layer becomes slightly different from its original version. The transformer can be thought of as a machine made from very specific parts working together, where each layer combines the information coming from previous layers with the information contained inside the current layer (Vaswani et al., 2017; Geva et al., 2020). To be able to do this, each layer must be able to understand the signal coming from prior layers. In simpler words, there is a notion of compatibility between the different layers of the transformer when they are trained together. As we edit one specific layer of the model continuously while keeping the rest of the model constant, we are constantly changing one part of the model while keeping the remaining part the same. Such a procedure is bound to reach a point where the layer that is changed becomes so different from its original version that this compatibility is destroyed and other parts of the transformer are unable to makes sense of the incoming signal from the layer being edited.
@@
Figure 3: This figure shows the editing proficiency of MEMIT on GPT-J for Sample 1 over 2000 sequential edits made to the model.
@@
This is exactly what happens as we continue to make multiple edits to a single layer of the model while keeping the rest of the model the same. This can be seen in Figure 1(d). Figure 1(d) shows the normalized [^5] We first take the L2 norm between original and post-edit weights of the layer being edited, and then normalize it by number of neurons in the layer. L2 distance between the weights of the edited layer and the original weights of the layer as a function the number of edits made to the model. We find that as more edits are made to the model, the distance between the original and edited layer continuously increases until it suddenly explodes. This is the point where the edited layer becomes incompatible with the rest of the model. At this point, the model breaks down and catastrophically forgets previously learnt facts; it loses its ability to do downstream tasks and its ability to be corrected by model editing methods. The gradual increase in distance between the original weight and new weights leads to the gradual region of forgetting, whereas the spike in the distance with a single updates leads to catastrophic forgetting.
3.4.1 Disabling Edits in ROME
@@
Finally, we take a deeper look at the specific edits that cause the inflection point in ROME. We call these edits disabling edits, as they disable the model and make it unusable for downstream tasks. Note that these are facts that ROME is successfully able to edit in the model. Are these disabling edits a result of continuous sequential editing that accumulates over time or of one specific fact that is especially hard for a model to learn?
We find that when we edit the facts corresponding to disabling edits as the first edit made to the model, the model is still left completely disabled, and the normalized distance of the layer weights from the original weights is comparable to the distances seen around the spikes. This can be seen in Table 2, where we present the average normalized L2 norm between the edited model layer and its original weights when only one edit is made to the model. We find that disabling edits have three orders of magnitude larger distance than non-disabling edits. This shows that the disabling edits in ROME are not a result of continuous sequential editing of the model, but a fundamental limitation of ROME. We can describe disabling edits as facts that ROME is unable to successfully edit without crippling the model. Such disabling edits can also be a source of potential adversarial attacks.
Property FT-C ROME MEMIT
EDITING EFFICACY 100% 100% until CF < 100% until CF
EDIT LOCALITY Very Low High Very High
AVERAGE DURATION BEFORE CF CF not observed Short Long
DOWNSTREAM PERFORMANCE LOSS High High Low
FACT FORGETTING PERCENTAGE Very High High Low
SINGLE DISABLING EDIT False True False
Table 3: Comparison between ROME and MEMIT at scale. CF refers to the point of catastrophic forgetting.
@@ 4 Scaling MEMIT
In this section, we evaluate the performance of the Mass-Editing Memory in a Transformer (MEMIT) method Meng et al. (2022b) when multiple sequential edits are made to the same model. We will follow the same procedure as followed in section 3, first evaluating the editing proficiency, fact forgetting and loss of performance on downstream tasks. We perform sequential editing on GPT-J (6B) using a random subset of 2000 examples from the CounterFact dataset when using MEMIT. This subset is a continuation of sample 1 in section 3.
@@
Figure 2(a) shows the editing proficiency of MEMIT as a function of the number of edits made to the model. Note that here we edit one fact at a time for MEMIT. While MEMIT is able to make batched edits, we leave that analysis for future work. The dotted lines show a window size of 5 previous edits, whereas solid lines show a window size of 50 previous edits, same as in Figure 0(c). We see that the efficacy score for MEMIT is not as high as ROME. This means that knowledge edits made via MEMIT are not always successful, while in ROME we’re always able to edit facts successfully. We also see a continuous decline of neighborhood score for MEMIT as seen for ROME, showing that editing facts also start affecting other facts stored in the model.
Figure 2(b) shows the percentage of successfully edited facts that get forgotten as new facts are edited using MEMIT. We again begin to see two phases of forgetting - gradual and catastrophic. The catastrophic forgetting phase begins after approximately 1400 edits made to the model. When compared to ROME, we find that edits made using MEMIT have a much longer gradual forgetting phase across multiple samples. Additionally, we also find that MEMIT forgets fewer previously edited facts when compared to ROME, as seen in Figure 4, where MEMIT forgets almost three times fewer facts when compared to ROME. This can also be seen for other samples in appendix A.5.2.
@@
Finally, the effect of the number of edits on the downstream performance of the model can be seen in Figure 2(c). We see that the model maintains its ability to downstream tasks as more number of edits are made to the model. While this is true for GPT-J, we observe a gradual loss of performance with multiple edits for GPT2-XL (appendix A.4). In fact, we find that for GPT2-XL, the model loses the ability to do paraphrase detection long before the point of catastrophic forgetting, thus showcasing that the model can lose its ability of performing certain downstream tasks even before a disabling edit cripples the model.
@@
Figure 2(d) shows the distance of the edited layers from the respective original layers. We find that the distance from the respective original layer increases gradually until approximately 1400 edits. After this, we find spikes in the distance between the edited layers and the original layers, which coincides with the points of catastrophic forgetting as seen in previous plots. We also evaluate if disabling edits are fundamental to MEMIT. We find that, if the fact that disables the model after a sequence of edits is edited first in the model, it does not lead to catastrophic forgetting. Thus, MEMIT is more robust to a single destabilizing edits when compared to ROME. We conjecture this is because a fact is stored within multiple layers of a model (Meng et al., 2022a, b), and editing the weights of a single layer to edit facts can lead to larger instabilities in the model. We summarize the properties of ROME and MEMIT at scale in Table 3, clearly showing MEMIT as a superior method across different parameters except editing efficacy.
@@ 5 Conclusion
In this paper we analyze popular model editing techniques at scale. We use these methods to make multiple sequential edits to the same model and find that they fail in multiple ways. We find that ROME and MEMIT perform the best when scaled to multiple sequential edits as measured using metrics like fact forgetting and downstream performance. As we edit the models, we discover they undergo forgetting of previous knowledge and skills in two phases. Initially, the model gradually forgets previously edited facts and loses the ability to do downstream tasks, a phase which we call gradual forgetting. After that, the model abruptly loses all coherence and function including the ability to recall previously edited facts, perform downstream tasks and the ability to be edited, which is a realization of catastrophic forgetting. We also find that the source of these two phases of forgetting is that the layers being edited with these methods slowly drift away from their original weight values, thus becoming incompatible with the rest of the model.
Figure 4: Compares the forgetting rate between ROME and MEMIT.
Practical use of model editing requires us to be able to make multiple sequential edits to a model. We find that these model editing methods, like other fine tuning techniques, are prone to catastrophic forgetting. To be able to scale such methods, we not only need to have high efficacy, specificity and generalization, but we also need these methods to preserve the model’s existing abilities. With this paper, we call for an improved evaluation of model editing techniques at scale, including evaluating model performance on downstream tasks and ability to recall previously edited facts.
Finally, we want to stress upon the implications of the two phases of forgetting discovered in this paper for ROME and MEMIT. Gradual forgetting makes model editing techniques increasingly less effective as we sequentially edit facts, and hence limits their usefulness at scale. While catastrophic forgetting, which renders the model practically useless, caps the extent to which we can scale these methods. Thus we need to create model editing techniques that can counteract both gradual forgetting and catastrophic forgetting when scaled.
6 Limitations
The aim of our work is to present the efficacy of current model editing techniques at scale and the usefulness of our proposed evaluation framework when studying model editing techniques at scale. To do so, in this paper we study models of size 1.5 billion and 6 billion parameters, which are standard models used in previous works (Mitchell et al., 2021; Meng et al., 2022a, b). While we see consistent behavior of all model editing methods for the two sizes, it is possible that as models grow even larger, they respond differently to different model editing techniques. Additionally, some model editing methods like MEND Mitchell et al. (2021) and MEMIT Meng et al. (2022b) have the ability to perform batched edits, that is, make multiple edits is one gradient update. Effects of model editing techniques on larger model sizes, batch edits with increasing batch sizes, as well combining multiple batches of edits sequentially are not presented in this paper. We find that these aspects of model editing are a natural extension of our work but were out of scope for this paper due to space constraints. Yet these settings can be easily evaluated under the framework we have presented in this paper and have been left for future work.